caira12:

simplify -3 1/9 - (-8 1/3)

Hero:

Hint: \(a \dfrac{b}{c} = a + \dfrac{b}{c}\) which means you can write the given expression as \(-3 + \dfrac{1}{9} - \left(-8 + \dfrac{1}{3}\right)\) The next step would be to distribute the negative across the parentheses to get \(-3 + \dfrac{1}{9} + 8 - \dfrac{1}{3}\) Afterwards you group like terms together: \((8 - 3) + \left(\dfrac{1}{9} - \dfrac{1}{3}\right)\) Then you finish simplifying from there.

Hero:

@caira12 can you finish from here?

caira12:

i think so

caira12:

so the final answer would be 5 2/9

Hero:

Actually, you should do this one step at a time. \((8 - 3) + \left(\dfrac{1}{9} - \dfrac{1}{3}\right)\) What is the fraction equivalent to 1/3 that gives it the same denominator as 1/9?

Hero:

Sorry you left so fast IDK why but I was trying to show you how to get the correct answer. You were on the right track. You just made a sign mistake. \((8 - 3) + \left(\dfrac{1}{9} - \dfrac{1}{3}\right)\) \(=5 + \left(\dfrac{1}{9} - \dfrac{3}{9}\right)\) \(=5 - \dfrac{2}{9}\) \(=-\left(5 + \dfrac{2}{9}\right)\) \(=-5\dfrac{2}{9}\) In other words, the resulting fraction is negative, not positive.